For the polynomial function f(x) = −5x(x + 2)2(x − 1)3:For the polynomial function f(x) = -5x(x + 2)(x - 1): 7. (4 points) The leading term when expanded is -5.2. Use this to des
Answer #1

7.

f(x) = -5x(x+2)^2(x-1)^3

leading \ coefficient -5x^6

so,

x\rightarrow \infty , f(x)\rightarrow -\infty

x\rightarrow -\infty , f(x)\rightarrow -\infty

8.

zero 0 -2 1
multiplicity 1 2 3

9.

for horizontal asymptote to be y = -5 .

lim_{x\rightarrow \infty }y = -5

so,

y = \frac{-5x}{x-2}